OpenStudy (akonkel):

What is the reciprocal of -3/4? @jagr2713

jagr2713 (jagr2713):

didnt dan give u the answer?

OpenStudy (akonkel):

ya but do u flip the negative to or no?

jagr2713 (jagr2713):

do you mean switch or flip

OpenStudy (akonkel):

switch

OpenStudy (anonymous):

-3/4 = 4/-3

OpenStudy (anonymous):

Either way, it'll be negative.

OpenStudy (anonymous):

Since: -3/4 = -(3/4) 4/-3 = -(4/3)